2014 Feb 23 (UT)

Jupiter Observations:

Corrected Pointing

Focus/Collimate --- last pupil shows evidence of higher order spherical. We had some trouble with FPIA not fitting pupils - it found pupils but did not fit these. MV was not running and was turned on.

MST 20:00 start J1, J2 ... J1 is a bit elongated.

Refocus/collimate again @ J2 field --- now focus is oscillating (+/- 1000 nm)

MST 20:22 start again with J1 (032108) seeing ~0.8" from tech chip,

MST 20:26 J2,

MST 20:29 J1 (033043), J2 20:36 (033533)

MST 20:39 J1 (034026), J2 (034522)

MST 20:50 J1 & J2 once more

Finished with LBC observations. There was quite a bit of cloud - roughly 1 magnitude of extinction.

Between MST 21:00 and ~00:30 TCS testing.

MST 00:45 reconfigured for MODS & LBCR. There is more cloud now - our 6-th magnitude guide star is going in/out of the clouds.

John is testing the pseudo-monocular script, taking 120-sec LBC-R images while on the pointing/collimation star. Script is in /home/lbcobs/supportscripts/SymbioticScripts/ and is pmm_observe.py. The command, ./pmm_observe.py -s DX -t 120, will take 120-sec exposures with LBC-Red so long as the telescope is "on-source". When one slews to a new field, the LBC-Red observations should resume, though the observer should stop to focus/collimate.

Slew to Feige 34: Had to choose a brighter guide star, as the R=14.12 one in the standard script was not seen... Chose guide star with R=10.99 (used PA= -90). The clouds were so thick and variable that the guide star appeared to be blinking on and off.

MST 01:55 Closed dome.

MST 02:45 Still thick clouds and dome remains shut. Running LBC-Blue (out of beam) and Red (in beam) with 60-sec exposures.

MST 02:51 Opening dome.

MST ~03:00 acqMODS feige67.acq. Run dofpia on redonly. Take mods1r.20140223.0004.fits and 0005.fits (better collimated but clouds came again).

modsAlign -r mods1.20140223.0005.fits to align and make offset: -3.363, 5.751 rel.

After offset, start .pmm_observe.py script to take red exposures... that is working. MODS 0006 is a confirmation image, but there is too much cloud to see star... seeing ~1.2".

execMODS feige67.obs --- very cloudy (noted in object title). For spectral response function only... but with the R=14 guide star, this is a struggle and it is not wave-front-sensing any longer. Will probably stop if it doesn't improve and move to the brightest quasar which is in the N, which seems to be the only place we see stars on the all-sky.

b 1 r 8 --- thick cloud b 2 r 9 --- better (seeing ~1"). AM=1.047 b 3 r 10 --- better but losing about 1 mag from previous exposure

MST 03:30 Slew

acqMODS J1435_1asec_HA-2to0.acq (PA=-30 is good for HA=-2:00 to -00:30).

pmm_observe.py canceled the OB, but LBC GUI says "command rejected since another is running!".
/home/lbcobs/supportscripts/SymbioticSupport/stop-ob.pl
38 Sun Feb 23 10:31:47 2014 MODS directGregorian LBC prime 0 0 | 1 0 STOPPED
LBC: STOPPED 0/0 g-SLOAN ready 0/0 r-SLOAN ready 0 0  
cancel OB for off-source on right
/home/lbcobs/supportscripts/SymbioticSupport/stop-ob.pl
39 Sun Feb 23 10:31:51 2014 MODS directGregorian LBC prime 0 0 | 1 0 STOPPED
LBC: STOPPED 0/0 g-SLOAN ready 0/0 r-SLOAN ready 0 0  
cancel OB for off-source on right
/home/lbcobs/supportscripts/SymbioticSupport/stop-ob.pl

Cntl-c'd the script. Unfortunately, the guide star (R=12.5) is not consistently "found" ...

Doesn't make sense to try dofpia to refocus/recollimate LBC-Red, but I restarted the script --- seems to be running fine.

MST 04:00 Nominally start of LBC-Red saturn observations but on the all-sky it looks completely overcast and there is not a good patch. Waiting...

MST 04:07 While waiting - fiddling with "pmm_observe", which is still running. Clicked "pause" on the LBC-GUI, pmm_observe restarted the observations. Clicked "stop on the GUI and again, pmm_observe resumed the observations, as expected. If there were stars bright enough to see on the AO tech chip, #1, these symbiotic LBC-Red observations could test the sending of "z4" corrections, but there are not.

MST 04:20 Watching clouds --- Noticed that offset from modsAlign had failed because no star was found when guiding had resumed. Took another exposure since still on-source and guide star is again seen. Offset had completed - just the error resulted because guiding could not be resumed.

MST 04:24 Clearing a little bit --- Will try the LBC targets, even though they are very faint and there is a lot of cloud. The cloud does seem very patchy.

MST 04:32 Slew to pointing field for Saturn observations to see how bad the transparency is here and to get ready in case it clears.

MST 04:40 Having trouble sending a preset with LBC. Deauthorizing MODS. OK now.

MST 04:47 No pupils found... Very cloudy...

MST 04:52 Closing dome. Putting MODS to sleep. Taking a series of 20 LBC-B and R images with no telescope or guiding, just to see about RPC/shutter issues. OB is /home/lbto/Calib_OBs/TCStesting/shuttertest.ob and uses the central region of chip 2 only.

MST 05:30 Took a series of 25 biases with dome shut & dome lights off. (Somehow just before this, I managed to click something on the GUI that caused LBC to go down - but it restarted).

-- %USERSIG{OlgaKuhn - 2014-02-23}%

Comments

 
Topic revision: r3 - 23 Feb 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback